A force of 30 N is applied to a body of mass 5.0 kg which is moving at speed 4.0 m/s. The final speed of the body after 2.0 a is [16] m/s
1
Expert's answer
2019-11-20T09:54:49-0500

Solution:

"F=m\times a"


"a=\frac{F}{m}"


"a=\frac{30 N}{5 kg}=6\frac{m}{s^2}"


"v=v_0+a\times t"

"t=\frac {v-v_0}{a}=\frac{16-4}{6}= 2 s"

Answer: 2s